# Permutahedron Triangulations via Total Linear Stability and the Dual Braid Group

### v1

@@ 1.1.1. $K(\pi,1)$ hyperplane complements
{paracol}
[…]
@@ 1.1.2. The Salvetti and Bessis–Brady–Watt complexes
{paracol}
[…]
@@ 1.1.3. The Artin and dual presentations
{paracol}
[…]
@@
[…]
These two different presentations for the braid group come from the two different combinatorial models for the complexified hyperplane complement mentioned in Section 1.1.2. In more detail: {paracol}2 For a finite Coxeter group $W$, the Salvetti complex ${\sf Sal}(W)$ can be taken to be invariant with respect to the action of the group, giving a homotopy equivalence between the quotient complex and the orbit space [71].
[…]
@@ 1.1.4. Presentations of the pure braid group
Extending the fiber-type technique from type $A$, Cohen gave a presentation of the pure braid group in type $B$ [25, Theorem 1.4.3] (see also Digne–Gomi [32, 31]). Markushevich claimed to have found a presentation in type $D$ in [52], but Falk and Randall later determined it was incorrect [35]. These approaches share similar problems with Artin’s presentation: they are type-specific and give complicated, unmotivated presentations. For example, Artin’s presentation uses five families of relations, while the presentations of Cohen and Digne–Gomi each use nine families.
@@ 1.3.4. The common geometric setting
$\Delta_{c}^{+}:=\bigcap_{t\in T_{\mathrm{L}}(c)}H_{t}^{-}=\bigcup_{u\in W_{c}^{+}}\overline{u^{-1}\mathbb{B}}=(1-c^{-1})^{-1}\mathrm{span}_{\geq 0}(\Phi^{+})^{\vee}=\subseteq V.$
(Here, $\mathrm{span}_{\geq 0}$ denotes nonnegative span.) For each $t\in T$, there is a distinguished vertex $\delta_{t}=(1-c^{-1})^{-1}\beta_{t}^{\vee}$} inside $\Delta_{c}^{+}$, and $\Delta_{c}^{+}=\mathrm{span}_{\geq 0}\{\delta_{s}:s\in S\}$. We will always draw $\Delta_{c}^{+}$ as a simplex by deleting the origin and passing to the positive projectivization (identifying points that have the same positive span).
@@
  • For $u\in W_{c}^{+}$, $A\in\mathrm{MClus}^{+}(W,c^{-1})$, and $\vec{\pi}\in{\sf MC}(W,c)$, each of the cones $\Delta(u)$, $\Delta(A)$, and $\Delta(\vec{\pi})$ is $r$-dimensional; it is the nonnegative span of some $r$-subset (determined from $u$, AA, or $\vec{\pi}$) of the vertex set $\{\delta_{t}:t\in T\}$.
@@
Continuing Example 1.11, we have redrawn $\Delta_{c}^{+}$ on the left side of Figure 5, showing the refinement of $\mathrm{MClus}^{+}(W,c^{-1})$ by $W_{c}^{+}$. Each simplicial cone is drawn as a triangle. Each distinguished vertex $\delta_{(i\,j)}$ (for $(i\,j)\in T$) is represented by a circle filled with $ij$. Let $u=s_{3}s_{2}\in W_{c}^{+}$. On the right side, we have marked $\Delta(u)$ with a white star (this triangle is labeled by $u^{-1}=s_{2}s_{3}$ on the left). Let $A\in\mathrm{MClus}^{+}(W,c^{-1})$ be the unique positive $c^{-1}$-cluster such that $u\in W_{c}^{+}(A)$. We have shaded the triangle $\Delta(A)$ in blue. The union of the blue and purple triangles is $\Delta(\vec{\pi})=\Delta(\vec{\pi}^{\prime})$, where $\mathrm{rw}(\vec{\pi})=(1\,3)(1\,2)(3\,4)$ and $\mathrm{rw}(\vec{\pi}^{\prime})=(1\,3)(3\,4)(1\,2)$.
@@ 1.3.5. Concordancy
Two medium triangles $M_{1}$ and $M_{2}$ corresponding to concordant pairs $(u_{1},\mathcal{C}_{1})$ and $(u_{2},\mathcal{C}_{2})$ are connected by an edge if and only there is an SBDW simplex in the collection $\{u_{1}\vec{\pi}\mathfrak{y}:\vec{\pi}\in\mathcal{C}_{1}\}$ that shares a facet with an SBDW simplex in the collection $\{u_{2}\vec{\pi}\mathfrak{y}:\vec{\pi}\in\mathcal{C}_{2}\}$.
@@ 5. Combinatorics of the SBDW Triangulation
We now prove the results from Section 1.3 in order to gain a much finer combinatorial understanding of SBDW triangulations. Recall that we write ${\sf Sort}(W,c)$ for the set of $c$-sortable elements of $W$, which are the minimum elements of the $c$-Cambrian congruence classes (in the right weak order).
@@ 6.1. Artin relations from dual relations
{paracol}
2 The Artin presentation of $\mathbf{B}_{\mathfrak{S}_{3}}$ is
@@
[…]
@@ 7.2. Noncrossing weak order and noncrossing Bruhat order
{paracol}
2 The positive Artin monoid $\mathbf{B}_{W,w_{\circ}}^{+}$ is the submonoid of $\mathbf{B}_{W,w_{\circ}}$ generated by $\mathbf{S}$; by Garside theory, this positive monoid injects into the braid group. The Coxeter group $W$ lifts to $\mathbf{B}_{W,w_{\circ}}^{+}$ by lifting each reduced $S$-word to the corresponding $\mathbf{S}$-word in Artin generators. We denote this positive Artin lift of $w\in W$ to $\mathbf{B}_{W,w_{\circ}}^{+}$ by $\mathbf{w}$.
[…]
@@ 7.4. Obstacles to an explicit homotopy
The issue in this $\mathfrak{S}_{3}$ example is that each of the simplices labeled by the $\preceq_{c}$-increasing chain are being “shared” by two permutahedral faces of ${\sf Sal}(\mathfrak{S}_{3})$.

### v2

@@ 1.1.1. $K(\pi,1)$ hyperplane complements
Building on Garside’s single mathematical paper [36] (“que je suis souvent de très près” [28]), Deligne proved that the complexified complement of a real finite simplicial hyperplane arrangement is a $K(\pi,1)$ space [28]. This result had previously been proven for the reflection arrangements of types $A,B,D,F_{4}$, and $I_{2}(m)$ by Brieskorn [21] (who also conjectured the result for all finite Coxeter groups). According to van der Lek [76], Arnold, Pham, and Thom subsequently conjectured it to hold for hyperplane arrangements for all (infinite) Coxeter groups [24, 58].
Building on the structure of Deligne’s proof [28] (“The general architecture of our proof is borrowed from Deligne’s original approach, but the details are quite different” [9]), Bessis proved that the complement of the hyperplane arrangement of a finite complex reflection group is a $K(\pi,1)$ space [9]. Various special cases [55, 56] (including those of Shephard groups, which have Artin groups isomorphic to those of certain real groups) were known prior to Bessis’s foundational work, and his “dual” approach has subsequently been generalized to affine Weyl groups [53] and Coxeter groups of rank $3$ [29].
@@ 1.1.2. The Salvetti and Bessis–Brady–Watt complexes
The complexified hyperplane complement of a real simplicial arrangement has a combinatorial model called the Salvetti complex. In [70], by choosing a real basepoint for the fundamental group, Salvetti used the real chamber geometry of any real finite hyperplane arrangement—not just simplicial arrangements—to construct a CW complex that is a deformation retract of the complexified hyperplane complement.
The hyperplane complement for a finite well-generated complex reflection group has a combinatorial model we call the Bessis–Brady–Watt complex. In [9], by choosing an eigenvector for a Coxeter element as the basepoint for the fundamental group, Bessis used the noncrossing partition lattice as a replacement for the chamber geometry for real arrangements (see also Brady and Brady–Watt [16, 17, 19]).
@@ 1.1.3. The Artin and dual presentations
The group $W$ is generated by its set $S$ of $r$ simple reflections, which act on $V$ via reflections through the walls of $\mathbb{B}$. The length $\ell_{S}(w)$ of an element $w\in W$ is the minimum number of simple reflections required to write $w$. Equivalently, $\ell_{S}(w)$ is the number of hyperplanes in $\mathcal{H}$ separating $\mathbb{B}$ from $w\mathbb{B}$.
The group $W$ is generated by its set $T$ of $N$ reflections, which act on $V$ via reflections through the hyperplanes in $\mathcal{H}$. The absolute length $\ell_{T}(w)$ of an element $w\in W$ is the minimum number of reflections required to write $w$. Equivalently (as $W$ is real), $\ell_{T}(w)$ is the dimension of the image of w1w-1.
@@
Writing $\bf c$} for the image of $c$ in $\mathbf{B}_{W,c}$, the full twist ${\bf c}^{h}$ lies in the center of $\mathbf{B}_{W,c}$.
These two different presentations for the braid group come from the two different combinatorial models for the complexified hyperplane complement mentioned in Section 1.1.2. In more detail:
For a finite Coxeter group $W$, the Salvetti complex ${\sf Sal}(W)$ can be taken to be invariant with respect to the action of the group, giving a homotopy equivalence between the quotient complex and the orbit space [71].
The Artin presentation for $\mathbf{B}_{W}$ is associated to this quotient Salvetti complex ${\sf Sal}(W)/W$: informally, this complex is formed by gluing together oriented copies of the $W$-permutahedron $\mathrm{Perm}_{\mathfrak{y}}(W)$—one for each element of WW—and then quotienting by the action of $W$. The edges of this complex can be viewed as oriented edges of the permutahedron, so they are labeled by simple reflections, while relations come from higher-dimensional faces of the permutahedron.
@@ 1.1.4. Presentations of the pure braid group
Extending the fiber-type technique from type $A$, Cohen gave a presentation of the pure braid group in type $B$ [25, Theorem 1.4.3] (see also Digne–Gomi [32, 31]). Markushevich claimed to have found a presentation in type $D$ in [52], but Falk and Randell later determined it was incorrect [35]. These approaches share similar problems with Artin’s presentation: they are type-specific and give complicated, unmotivated presentations. For example, Artin’s presentation uses five families of relations, while the presentations of Cohen and Digne–Gomi each use nine families.
@@ 1.3.4. The common geometric setting
$\Delta_{c}^{+}:=\bigcap_{t\in T_{\mathrm{L}}(c)}H_{t}^{-}=\bigcup_{u\in W_{c}^{+}}\overline{u^{-1}\mathbb{B}}=(1-c^{-1})^{-1}\mathrm{span}_{\geq 0}(\Phi^{+})^{\vee}\subseteq V.$
(Here, $\mathrm{span}_{\geq 0}$ denotes nonnegative span.) For each $t\in T$, there is a distinguished point $\delta_{t}=(1-c^{-1})^{-1}\beta_{t}^{\vee}$} inside $\Delta_{c}^{+}$, and $\Delta_{c}^{+}=\mathrm{span}_{\geq 0}\{\delta_{s}:s\in S\}$. We will always draw $\Delta_{c}^{+}$ as a simplex by deleting the origin and passing to the positive projectivization (identifying points that have the same positive span). We will sometimes refer to the points $\delta_{t}$ for $t\in T$ as vertices since they will naturally correspond to vertices of the positive $c^{-1}$-cluster complex.
@@
  • For $u\in W_{c}^{+}$, $A\in\mathrm{MClus}^{+}(W,c^{-1})$, and $\vec{\pi}\in{\sf MC}(W,c)$, each of the cones $\Delta(u)$, $\Delta(A)$, and $\Delta(\vec{\pi})$ is $r$-dimensional; it is the nonnegative span of some $r$-subset (determined from $u$, AA, or $\vec{\pi}$) of the set $\{\delta_{t}:t\in T\}$.
@@
Continuing Example 1.11, we have redrawn $\Delta_{c}^{+}$ on the left side of Figure 5, showing the refinement of $\mathrm{MClus}^{+}(W,c^{-1})$ by $W_{c}^{+}$. Each simplicial cone is drawn as a triangle. Each distinguished point $\delta_{(i\,j)}$ (for $(i\,j)\in T$) is represented by a circle filled with $ij$. Let $u=s_{3}s_{2}\in W_{c}^{+}$. On the right side, we have marked $\Delta(u)$ with a white star (this triangle is labeled by $u^{-1}=s_{2}s_{3}$ on the left). Let $A\in\mathrm{MClus}^{+}(W,c^{-1})$ be the unique positive $c^{-1}$-cluster such that $u\in W_{c}^{+}(A)$. We have shaded the triangle $\Delta(A)$ in blue. The union of the blue and purple triangles is $\Delta(\vec{\pi})=\Delta(\vec{\pi}^{\prime})$, where $\mathrm{rw}(\vec{\pi})=(1\,3)(1\,2)(3\,4)$ and $\mathrm{rw}(\vec{\pi}^{\prime})=(1\,3)(3\,4)(1\,2)$.
@@ 1.3.5. Concordancy
Two medium triangles $M_{1}$ and $M_{2}$ corresponding to concordant pairs $(u_{1},\mathcal{C}_{1})$ and $(u_{2},\mathcal{C}_{2})$ are connected by an edge if and only if there is an SBDW simplex in the collection $\{u_{1}\vec{\pi}\mathfrak{y}:\vec{\pi}\in\mathcal{C}_{1}\}$ that shares a facet with an SBDW simplex in the collection $\{u_{2}\vec{\pi}\mathfrak{y}:\vec{\pi}\in\mathcal{C}_{2}\}$.
@@ 5. Combinatorics of the SBDW Triangulation
We now prove the results from Section 1.3 in order to gain a much finer combinatorial understanding of SBDW triangulations, whose simplices are labeled by concordant pairs $(u,\vec{\pi})\in W_{c}^{+}\times{\sf MC}(W,c)$. We characterize when elements of $W_{c}^{+}$ are concordant with the same set of chains in ${\sf MC}(W,c)$ using Cambrian congruence classes, show the subset of $W_{c}^{+}$ concordant with a fixed $\vec{\pi}\in{\sf MC}(W,c)$ is an interval in the left weak order, and characterize when $u\in W_{c}^{+}$ is concordant with $\vec{\pi}\in{\sf MC}(W,c)$ using the positive cluster complex.
Recall that we write ${\sf Sort}(W,c)$ for the set of $c$-sortable elements of $W$, which are the minimum elements of the $c$-Cambrian congruence classes (in the right weak order).
@@
We will need the next lemma in what follows.
@@ 6.1. Artin relations from dual relations
The Artin presentation of $\mathbf{B}_{\mathfrak{S}_{3}}$ is
@@
One can interpret the expression $\color[rgb]{0,0.7,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.7,1}\mathbf{s}_{1}\mathbf{s}_{2}\mathbf{s}_{1}$} as the result of walking from $\mathfrak{y}$ to $w_{\circ}\mathfrak{y}$ along the left-hand edges of the $\mathfrak{S}_{3}$-permutahedron $\mathrm{Perm}_{\mathfrak{y}}$ in Figure 1. Similarly, the expression $\color[rgb]{1,0.7,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.7,0}\mathbf{s}_{2}\mathbf{s}_{1}\mathbf{s}_{2}$} is the result of walking from $\mathfrak{y}$ to $w_{\circ}\mathfrak{y}$ along the right-hand edges of $\mathrm{Perm}_{\mathfrak{y}}$. The equality of these two expressions in the braid group corresponds to a $2$-cell, which is all of $\mathrm{Perm}_{\mathfrak{y}}$.
Bessis’s dual presentation of $\mathbf{B}_{\mathfrak{S}_{3}}$ is
@@ 7.2. Noncrossing weak order and noncrossing Bruhat order
The positive Artin monoid $\mathbf{B}_{W,w_{\circ}}^{+}$ is the submonoid of $\mathbf{B}_{W,w_{\circ}}$ generated by $\mathbf{S}$; by Garside theory, this positive monoid injects into the braid group. The Coxeter group $W$ lifts to $\mathbf{B}_{W,w_{\circ}}^{+}$ by lifting each reduced $S$-word to the corresponding $\mathbf{S}$-word in Artin generators. We denote this positive Artin lift of $w\in W$ to $\mathbf{B}_{W,w_{\circ}}^{+}$ by $\mathbf{w}$.
The map $w\mapsto\mathbf{w}$ is an isomorphism from the right weak order on $W$ to the interval $\mathbf{e},\mathbf{w}_{\circ}]_{\mathbf{S}}$ in $\mathbf{B}_{W,w_{\circ}}^{+}$.
The dual braid monoid $\mathbf{B}_{W,c}^{+}$ is the submonoid of $\mathbf{B}_{W,c}$ generated by $\mathbf{T}_{c}$; by Garside theory, this positive monoid injects into the braid group. The $c$-noncrossing partitions lift to $\mathbf{B}_{W,c}^{+}$ by lifting each reduced $T$-word to the corresponding $\mathbf{T}_{c}$-word in dual generators. We denote this positive dual lift of $\pi\in{\sf NC}(W,c)$ to $\mathbf{B}_{W,c}^{+}$ by $\bm{\pi}_{c}$.
@@ 7.4. Obstacles to an explicit homotopy
The issue in this $\mathfrak{S}_{3}$ example is that each of the simplices labeled by the $\preceq_{c}$-increasing chain appears in two permutahedral faces of ${\sf Sal}(\mathfrak{S}_{3})$.
@@ Acknowledgments
Appendix A Upgrading triangulations to subdivisions
Our main result in this subsection is the following proposition, which allows us to continuously deform subdivisions under certain conditions.
Proposition A.1.
Let $(P_{\tau})_{\tau\in[0,1]}$ be a collection of $d$-dimensional polytopes in $U$. Let $\mathfrak{J}_{1},\ldots,\mathfrak{J}_{m}\subseteq\mathfrak{V}(P_{0})$. Suppose that for each $\tau\in[0,1]$, there is a bijection $\varphi_{\tau}\colon\mathfrak{V}(P_{0})\to\mathfrak{V}(P_{\tau})$. Define $Q_{\tau}^{i}:=\mathrm{conv}(\varphi_{\tau}(\mathfrak{J}_{i}))$. Assume the following conditions hold:
  • $\{Q_{0}^{i}\}_{i=1}^{m}$ is a subdivision of $P_{0}$
  • The map $\varphi_{0}$ is the identity map on $\mathfrak{V}(P_{0})$.
  • For every $\tau\in[0,1]$, there is a poset isomorphism from $\mathsf{Faces}(P_{0})$ to $\mathsf{Faces}(P_{\tau})$ that agrees with $\varphi_{\tau}$ on $\mathfrak{V}(P_{0})$.
  • For every $\tau\in[0,1]$, there is a poset isomorphism from $\mathsf{Faces}(Q_{0}^{i})$ to $\mathsf{Faces}(Q_{\tau}^{i})$ that agrees with $\varphi_{\tau}$ on $\mathfrak{V}(Q_{0}^{i})=\mathfrak{J}_{i}$.
Then $\{Q_{\tau}^{i}\}_{i=1}^{m}$ is a subdivision of $P_{\tau}$ for every $\tau\in[0,1]$.
Throughout this subsection, we preserve the hypotheses and notation from Proposition A.1. Before proving that proposition, we need the following lemmas.
Lemma A.2.
Suppose $i,i^{\prime}\in[m]$ are such that $Q_{0}^{i}$ and $Q_{0}^{i^{\prime}}$ intersect along a common facet, which is necessarily $\mathrm{conv}(\mathfrak{J}_{i}\cap\mathfrak{J}_{i^{\prime}})$. For every $\tau\in[0,1]$, the polytopes $Q_{\tau}^{i}$ and $Q_{\tau}^{i^{\prime}}$ intersect along a common facet, which is $\mathrm{conv}(\varphi_{\tau}(\mathfrak{J}_{i}\cap\mathfrak{J}_{i^{\prime}}))$.
Proof.
Let $\mathfrak{w}\in\mathfrak{J}_{i}\setminus\mathfrak{J}_{i^{\prime}}$ and $\mathfrak{w}^{\prime}\in\mathfrak{J}_{i^{\prime}}\setminus\mathfrak{J}_{i}$. We note that by assumption, $\mathrm{conv}(\varphi_{\tau}(\mathfrak{J}_{i}\cap\mathfrak{J}_{i^{\prime}}))$ is a facet of $Q_{\tau}^{i}$, and so in particular the affine hull of $\varphi_{\tau}(\mathfrak{J}_{i}\cap\mathfrak{J}_{i^{\prime}})$ is a hyperplane $H_{\tau}$. Let $\mathcal{T}$ be the set of $\tau\in[0,1]$ such that $\varphi_{\tau}(\mathfrak{w})$ and $\varphi_{\tau}(\mathfrak{w}^{\prime})$ lie in opposite open half-spaces bounded by $H_{\tau}$. Because each map $\tau\mapsto\varphi_{\tau}(\mathfrak{v})$ for $\mathfrak{v}\in\mathfrak{V}(P_{0})$ is continuous, $\mathcal{T}$ is an open subset of $[0,1]$ that contains 0. We wish to show that $\mathcal{T}=[0,1]$, so suppose by way of contradiction that this is not that case. Then $0,1]\setminus\mathcal{T}$ contains its minimum, which is some number $\tau^{\#}$. Either $\varphi_{\tau^{\#}}(\mathfrak{w})$ or $\varphi_{\tau^{\#}}(\mathfrak{w}^{\prime})$ lies in the hyperplane $H_{\tau^{\#}}$. This contradicts the assumption that $Q_{0}^{i}$ and $Q_{\tau}^{i}$ (or $Q_{0}^{i^{\prime}}$ and $Q_{\tau}^{i^{\prime}}$) are combinatorially isomorphic. We conclude that $\mathcal{T}=[0,1]$, as desired.
This shows that for all $\tau$, all elements of $\varphi_{\tau}(\mathfrak{J}_{i}\setminus\mathfrak{J}_{i^{\prime}})$ lie in the opposite open half-space defined by $H_{\tau}$ from all elements of $\varphi_{\tau}(\mathfrak{J}_{i^{\prime}}\setminus\mathfrak{J}_{i})$. So the intersection $Q_{\tau}^{i}\cap Q_{\tau}^{i^{\prime}}$ is contained in $H_{\tau}$. On the other hand, $Q_{\tau}^{i}\cap H_{\tau}$ and $Q_{\tau}^{i^{\prime}}\cap H_{\tau}$ are exactly the facet $\mathrm{conv}(\varphi_{\tau}(\mathfrak{J}_{i}\cap\mathfrak{J}_{i^{\prime}}))$. ∎
Lemma A.3.
Choose $i\in[m]$, $\tau\in[0,1]$, and $F$ a facet of $Q_{\tau}^{i}$. Either $F$ is contained in the boundary of $P_{\tau}$, or $F=Q_{\tau}^{i}\cap Q_{\tau}^{j}$ for some $j\neq i$. If Proposition A.1 holds for (d1)(d-1)-dimensional polytopes and $F$ is contained in the boundary of $P_{\tau}$, then for every $j\neq i$, the relative interior of $F$ does not intersect $Q_{\tau}^{j}$.
Proof.
The conclusion of the lemma certainly holds if $\tau=0$.
Now suppose $\tau>0$. By the fourth bulleted assumption of Proposition A.1, there is a facet $F_{0}$ of $Q_{0}^{i}$ such that $\mathfrak{V}(F)=\varphi_{\tau}(\mathfrak{V}(F_{0}))$. If $F_{0}$ is not contained in the boundary of $P_{0}$, then no facet of $P_{0}$ contains $\mathfrak{V}(F_{0})$. The third bulleted item of the hypotheses of Proposition A.1 imply that no facet of $P_{\tau}$ contains $\mathfrak{V}(F)$, so $F$ is not contained in the boundary of $P_{\tau}$ either. Also, since $\{Q_{0}^{a}\}_{a=1}^{m}$ is a subdivision, $F_{0}=Q_{0}^{i}\cap Q_{0}^{j}$ for some $j\neq i$. By Lemma A.2, $Q_{\tau}^{i}\cap Q_{\tau}^{j}=\mathrm{conv}(\varphi_{\tau}(\mathfrak{J}_{i}\cap\mathfrak{J}_{j}))=F$ as desired.
If $F_{0}$ is contained in the boundary of $P_{0}$, then some facet of $P_{\tau}$ contains $\mathfrak{V}(F_{0})$. The third bulleted assumption of Proposition A.1 then implies that $F$ is contained in some facet of $P_{\tau}$.
For the final statement, suppose Proposition A.1 holds for (d1)(d-1)-dimensional polytopes. Choose a facet $R_{0}$ of $P_{0}$, and let $R_{\tau}$ be the facet of $P_{\tau}$ with $\mathfrak{V}(R_{\tau})=\varphi_{\tau}(\mathfrak{V}(R))$. Note that $R_{0}$ and $R_{\tau}$ are combinatorially isomorphic by assumption. Let $H_{\tau}$ be the corresponding facet hyperplane. Let $I$ be the set of $i\in[m]$ such that $Q_{0}^{i}\cap H_{0}$ is a (d1)(d-1)-dimensional polytope. Then $\{Q_{0}^{i}\cap H_{0}\}_{i\in I}$ is a subdivision of $R_{0}$. The deformations $\{Q_{\tau}^{i}\cap H_{\tau}\}_{i\in I}$ satisfy the assumptions of Proposition A.1 using the same map $\varphi_{\tau}$ (restricted to $\mathfrak{V}(R_{0})$). Thus, $\{Q_{\tau}^{i}\cap H_{\tau}\}_{i\in I}$ is a subdivision of $R_{\tau}$.
Suppose that $F$ is contained in $R_{\tau}$ and $F\cap Q_{\tau}^{j}\neq\emptyset$ for some $j\neq i$. This means $Q_{\tau}^{j}\cap H_{\tau}$ is nonempty and, since $H_{\tau}$ is a supporting hyperplane of $P_{\tau}\supset Q_{\tau}^{j}$, the polytope $Q_{\tau}^{j}\cap H_{\tau}$ is some face $F^{\prime}$ of $Q_{\tau}^{j}$; $\mathfrak{V}(F^{\prime})$ is exactly $\varphi_{\tau}(\mathfrak{J}_{j})\cap\mathfrak{V}(R_{\tau})$. By the fourth bulleted assumption of Proposition A.1, there is a face $F_{0}^{\prime}$ of $Q_{0}^{j}$ whose vertices are $\varphi_{\tau}^{-1}(\mathfrak{V}(F^{\prime}))=\mathfrak{J}_{j}\cap R_{0}$. We have $F_{0}^{\prime}=Q_{0}^{j}\cap H_{0}$. Since $\{Q_{0}^{i}\}_{i=1}^{m}$ is a subdivision, $F_{0}$ and $F_{0}^{\prime}$ intersect in a common face contained in $H_{0}$. Since $\{Q_{\tau}^{i}\cap H_{\tau}\}_{i=1}^{m}$ is a subdivision, $F$ and $F^{\prime}$ also intersect in a common face. Thus, the relative interior of $F$ does not intersect $Q_{\tau}^{j}$.
Proof of Proposition A.1.
We proceed by induction on dimension. The base case of $d=0$ is trivial. Now assume the proposition holds for dimension d1d-1.
Let $\mathcal{T}$ be the set of $\tau\in[0,1]$ such that $\{Q_{\tau}^{i}\}_{i=1}^{m}$ is not a subdivision of $P_{\tau}$; our goal is to show that $\mathcal{T}=\emptyset$. It follows from Lemma 3.1 and the hypothesis of the proposition that $\bigcup_{i=1}^{m}Q_{\tau}^{i}=P_{\tau}$ for every $\tau\in[0,1]$. Therefore, $\mathcal{T}$ is equal to the set of $\tau\in[0,1]$ such that two simplices in $\Theta_{\tau}$ do not intersect along a common face.
Let $\mathcal{T}^{\prime}$ be the set of $\tau\in[0,1]$ such that there exist $Q_{\tau}^{i},Q_{\tau}^{i^{\prime}}$ that intersect in their interiors. Since interiors of polytopes are open sets, we know that $\mathcal{T}^{\prime}$ is an open subset of $[0,1]$. We will show that $\mathcal{T}=\mathcal{T}^{\prime}$.
We certainly have $\mathcal{T}^{\prime}\subseteq\mathcal{T}$. Suppose that there exists $\tau\in\mathcal{T}\setminus\mathcal{T}^{\prime}$; we will obtain a contradiction, which implies the desired equality. According to Lemma 2.21, there exist $i,i^{\prime}\in[m]$ such that $Q_{\tau}^{i}\cap Q_{\tau}^{i^{\prime}}$ is (d1)(d-1)-dimensional but is not a common facet of $Q_{\tau}^{i}$ and $Q_{\tau}^{i^{\prime}}$. Then $Q_{\tau}^{i}\cap Q_{\tau}^{i^{\prime}}$ is properly contained in, and intersects the relative interior of, some facet $F$ of $Q_{\tau}^{i}$. Note that $F$ is not contained in the boundary of $P_{\tau}$. By Lemma A.3, there exists $j\in[m]$ such that $F=Q_{\tau}^{i}\cap Q_{\tau}^{j}$. But then $Q_{\tau}^{i^{\prime}}$ and $Q_{\tau}^{j}$ must intersect in their interior, which contradicts the assumption that $\tau\not\in\mathcal{T}^{\prime}$.
Now we would like to show that $\mathcal{T}^{\prime}$ is empty, which would prove the proposition. Fix $\tau\in[0,1]$, and suppose for the sake of contradiction that some point $\Pi$ is in the interior of both $Q_{\tau}^{i}$ and $Q_{\tau}^{j}$. Let $X$ be the union of the (d2)(d-2)-dimensional faces of the polytopes in $\Theta=\{Q_{\tau}^{i}\}_{i=1}^{m}$. Choose a point $\Pi^{\prime\prime}$ in $U\setminus P_{\tau}$ generically, so that the line segment from $\Pi$ to $\Pi^{\prime\prime}$ does not intersect $X$, and let $\Pi^{\prime}$ be the intersection of this line segment with the boundary of $P_{\tau}$. We claim that every point of the line segment $L$ from $\Pi$ to $\Pi^{\prime}$ is contained in at least two polytopes in $\Theta$. This claim leads to a contradiction: if $\Pi^{\prime}\in Q_{\tau}^{a}\cap Q_{\tau}^{b}$, then since $\Pi^{\prime}$ is in the boundary of $P_{\tau}$ and is not in $X$, it is in the relative interior of a facet of $Q_{\tau}^{a}$ and a facet of $Q_{\tau}^{b}$. But this contradicts Lemma A.3, as desired.
We now prove the claim. As you walk along $L$, you are contained in $Q_{\tau}^{i}\cap Q_{\tau}^{j}$ up until you arrive at a point $p$ where $L$ hits a facet $F$ of one of the polytopes, say $Q_{\tau}^{i}$, transversely. If $p=\Pi^{\prime}$, we are done, so we assume $F$ is not on the boundary of $P_{\tau}$. Lemma A.3 guarantees that $F=Q_{\tau}^{i}\cap Q_{\tau}^{a}$ for some $a$. If somehow $p$ is also the point where $L$ hits a facet $F^{\prime}$ of $Q_{\tau}^{j}$ transversely, again Lemma A.3 guarantees $F^{\prime}=Q_{\tau}^{j}\cap Q_{\tau}^{b}$ for some $b$. It is impossible for $F$ to equal $F^{\prime}$: in that case, the vertex sets agree, so the corresponding facets of $Q_{0}^{i}$ and $Q_{0}^{j}$ are equal, and since we have a subdivision for $\tau=0$, their intersection $Q_{0}^{i}\cap Q_{0}^{j}$ is their shared facet and Lemma A.2 implies that $Q_{\tau}^{i}\cap Q_{\tau}^{j}$ is a common facet, which is impossible as their interiors intersect. It is also impossible to have $a=b$ because $p\notin X$, so $p$ is in at most one facet of any polytope. In summary, if you go slightly past $p$ on $L$, you are still in two polytopes in $\Theta$. Repeating this argument as you proceed along $L$ proves the claim. ∎